# MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations

### v1

@@ 1 Introduction
Advances in large vision–language models (LVLMs) have significantly improved AI’s ability to interpret images and generate natural-language responses. However, existing benchmarks predominantly focus on short-form visual question answering [68, 30, 12, 51], captioning [53, 50], or grounded generation under constrained contexts [29]. These tasks fall short of capturing the interactive, decision-centric nature of real-world expert consultations, where users often present open-ended, ambiguous, and visually grounded queries. In knowledge-intensive domains such as agriculture, medicine, and engineering, expert consultations inherently span multiple modalities [49] and success hinges not just on perception or language fluency, but on the ability to reason causally, handle missing context, and make interaction-level decisions. This complex interplay between multimodal understanding and contextual reasoning in professional settings represents a significant gap in the current LVLM evaluation frameworks. LLMs and LVLMs are increasingly used in domains like healthcare, law, and plant care [57, 59, 38, 71, 17, 48, 63], yet current benchmarks underrepresent knowledge-intensive scenarios involving ambiguous, multimodal queries [30]. Agriculture illustrates this gap: farmers often seek image-based, expert-level guidance during critical events, but inaccurate or unsupported model outputs [69, 40, 67] can lead to serious consequences. This highlights the need for rigorous evaluation frameworks that assess VLMs in more realistic settings.
@@ 2.1 Overview of MIRAGE
MIRAGE comprises two components: MMST, with over 29,000 single-turn expert-user interactions, and MMMT, a multi-turn corpus of 6,306 dialogues ( 12,000 turns) requiring clarify-or-respond decisions. Tasks span identification, diagnostic reasoning, and issue management guidance across seven agronomic categories (Table 4). MMST includes Standard and Contextual subsets (Section 3, Appendix C), and a fine-grained evaluation framework assesses both accuracy and utility. With over 7,600 biological entities, including thousands of unseen species in the evaluation set, MIRAGE introduces a challenging open-world generalization setting for LVLMs. Refer Appendix C for more details.
@@
| iNat21 [58] | CLS | ✗ | ✓ | ✓ | – | – | – |
| TreeOfLife [54] | CLS | ✓ | ✓ | ✓ | – | – | – |
| SimpleQA [61] | OEQ | ✗ | ✗ | ✗ | ✓ | – | – |
| MMMU [68] | MCQ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ |
| AgMMU [30] | MCQ+OEQ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| CROP [70] | LFQ/Conv | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@ 2.2 Benchmark Comparison and Positioning
MIRAGE builds on a growing body of benchmarks in agriculture and multimodal reasoning, yet fills a critical gap left by prior efforts. Datasets like MMMU [68] focus on generalist multimodal reasoning but rely on constrained multiple-choice formats. AgMMU [30] targets agriculture, yet uses synthetic short-form MCQs that are not representative of real expert-user interactions. Domain-specific datasets such as TreeOfLife [54] and iNat21 [58] emphasize fine-grained image classification but operate in closed-world, non-interactive settings. The CROP benchmark [70] introduces multi-turn crop science QA, but is limited to two crops, lacks visual input, and is fully text-based. In contrast, MIRAGE is the only benchmark constructed from large-scale, real-world agricultural consultations and supports both multimodal single-turn and multi-turn tasks grounded in naturally underspecified user queries. As summarized in Table 1, MIRAGE includes contextual question answering, real user-submitted images with varied quality and lighting, and domain-expert-authored responses as ground truth. It spans diverse query types—including identification, causal reasoning, and management—and introduces goal-state modeling, clarify-or-respond decision policies, and a multidimensional evaluation framework that moves beyond correctness to assess identification accuracy, causal justification, response quality, and diagnostic parsimony. These features make MIRAGE a uniquely realistic and challenging testbed for evaluating LVLMs in real-world consultation scenarios.
@@ 3 MIRAGE-MMST: Multimodal Singleturn Benchmark
MIRAGE-MMST is a benchmark designed to assess expert-level, single-turn reasoning in multimodal agricultural consultations. The task setup is similar to a Long-form VQA task [37]. Each instance consists of a natural language question paired with one or more user-provided images and associated metadata (e.g., timestamp, location). Each instance consists of a natural language question , an associated image set , and metadata . Formally, a single-turn instance is represented as a triplet , and the model must generate a structured response , where denotes identified entities (e.g., crop, pest, disease), is a causal explanation, and is a management recommendation, if requested. The task evaluates the model’s ability to reason causally about visual symptoms, identify relevant agronomic entities, and, when prompted, generate detailed management recommendations grounded in the observed evidence.
To support varying levels of difficulty and contextual grounding, MIRAGE-MMST is divided into two subsets: a Standard subset, consisting of self-contained questions that can be answered using only the provided text and image, and a Contextual subset, where successful interpretation depends on implicit information such as time, location, or agricultural context not present in the input. The standard subset evaluates a model’s ability to identify entities, infer causal relationships, and generate explanatory or recommendation-based responses. In contrast, the contextual subset contains queries with latent information gaps and elliptical language, requiring models to reconstruct missing context using external priors. We first manually annotated a seed set of contextual examples and then adopted an automated classifier to separate the full dataset. See Appendix C.3 for details.
@@ 4.1 Dataset Curation
The goal of MIRAGE-MMMT is to model the decision-making process of an expert who must decide, based on the evolving multimodal context, whether to respond with actionable guidance or clarify by seeking more information. To simulate this, we extract segments of each dialogue such that the input context ends with a user utterance, and a later user message, typically written in response to an earlier expert query, is treated as a revealed fact. This allows us to reconstruct the decision point at which the expert, based on available context and images, would determine whether the current information is sufficient to respond or whether a clarification is necessary. We prompt a powerful LLM (e.g., GPT-4o) to generate a structured task instance consisting of dialogue context, referenced images, and metadata such as geographic location and topic. To ensure data safety, we perform automated PII sanitization while preserving domain relevance. (see Appendix F.2 for more details)
@@ 5.1 Model Selection
The LVLMs we evaluated can be divided into two groups: 1.) Proprietary Models: this group comprises SOTA GPT models (4.1, 4.1-mini, 4o, 4o-mini) [7], Claude-3 (3.5v2 Sonnet, 3.7 Sonnet, 3 Haiku) [11] available through API service. 2.) Open-Source Models: this group includes LLaMa-4 Scout-17B [44], LLaVa (7B mistral, 7B qwen2) [41], Qwen 2.5 VL models (3B, 7B, 32B, 72B) [5], Gemma 3 models (4b, 12, 27b) [56] and Intern VL3 (2B, 8B, 14B, 38B, 72B) [73]. 3.) Finetuned Models We perform multi-image multimodal fine-tuning of the Qwen 2.5 VL models on the MMST dataset. All experiments were conducted with NVIDIA A100-40GB/H200-141GB GPUs.
@@ 5.2 Evaluation Methodology
To facilitate consistent, interpretable, and reproducible evaluation of model outputs, we implement a structured LLM-as-a-Judge protocol. Rather than relying on single-model, single-pass assessments, we leverage an ensemble of reasoning-capable language models: DeepSeek-R1-Distilled [34], Qwen3-32B [5], and Phi-4-Reasoning [6], as judges. Each candidate’s response is evaluated across three independent generations, with every generation scored by the full ensemble, resulting in nine total evaluations per sample. This design captures both linguistic variability across generations and scoring consistency across evaluators. To validate the fidelity of this framework, we quantify inter-rater agreement using Fleiss’ for categorical scoring and Kendall’s W for ordinal coherence, ensuring both statistical reliability and scoring consistency across the ensemble. Fleiss’ scores for the ID task (binary classification) generally fall within the 0.75–0.88 range, indicating “good” to “excellent” agreement by standard interpretation guidelines. Full details, scoring rubrics, and agreement statistics are provided in Appendix D.1.1. The prompt for evaluation is showed in Figure 14.
@@ 5.3 MIRAGE-MMST Results
To evaluate model outputs on the MMST-MG task, we adopt a multidimensional scoring framework that assesses both factual correctness and communication quality. Each model response is rated on a 0–4 scale across the following four dimensions the judges: 1.) Accuracy: which measures the factual alignment with the expert’s response. A perfect score (4) is awarded when the facts, terminologies, and recommendations fully align with the expert response. Relevance measures the ability of the model to stay on-topic with the user’s original query and avoid tangential information. Completeness assesses whether the model covers all key information provided by the expert, and Parsimony, inspired by Occam’s Razor; this score captures the conciseness, clarity, and actionability of the model response. We aggregate these scores in a 2:1:1:1 ratio. See Appendix D.1.2 for more details.
@@ 5.4 Performance Comparison
| Acc (%) | Reason | Acc | Rel | Comp | Pars | W-Sum | |
@@
Scores are averaged over the three judge models. Closed-source rows are red/light-red, open-source rows blue/light-blue. Bold purple numbers mark the overall best for each metric.
@@
Model Finetuning: We utilize the MMST training dataset to fine-tune Qwen 2.5 VL models (3B, 7B, 32B) and evaluate performance on both MMST-ID & MMST-MG tasks. Figure 1 illustrates the effects of LoRA-based fine-tuning on the Qwen2.5-VL-3B model across five checkpoints. Our results indicate consistent improvements in seen entity accuracy on MMST-ID after fine-tuning, rising from 22.3% (base-Instruct) to a peak of 28.4% (epoch 6). Across model sizes, as seen in Figure 10, we see a consistent convergence behavior across epochs, with 32B achieving the strongest peak performance (37.6% at epoch 6) and reasoning score (3.04). However, models struggle to generalize, as accuracy on unseen entities remains stagnant. On the MMST-MG task, from the base-Instruct model to LoRA-ep6, we observe steady improvement across all four metrics as well. Relevance is the highest-scoring metric throughout, indicating the model’s ability to remain on-topic is strong even in zero-shot, and improves further with tuning. Parsimony also improves consistently, indicating models can learn concise delivery of recommendations.
@@ 5.6 Performance Comparison
We evaluate the model’s high-level decision and its alignment with user goals in three complementary ways. First, Decision Accuracy captures how often the model’s chosen action, whether to ask a clarification question or to respond directly, matches the gold annotation; To ensure that when the model generates an utterance, it is doing so meaningfully, we measure Clarify Relevance: for every turn labeled <Clarify> or Respond Relevance for all turns labeled <Respond> using the LLM judges.
@@
We conduct experiments under Zero-Shot and COT [60] prompt settings on MMMT shown in Table 3. In the zero-shot setting, GPT-4o achieves the highest Decision Accuracy (62.98%) and leads in both Clarify Goal Relevance (70.75%) and Respond Goal Relevance (77.07%), showing strong overall reasoning and generation alignment. Open models like LLaMA 4 Maverick and LLaMA 4 Scout perform competitively, reaching 53–49% decision accuracy. Introducing chain-of-thought (CoT) reasoning consistently improves performance across models. GPT-4o gains +2.5% in Decision Accuracy (to 65.5%) and improves goal relevance in both Clarify (+2%) and Respond (+1.5%) settings. Other models, especially Claude 3.7 and Qwen 72B, benefit more substantially (improvements in clarify goal relevance by +10.5%).
@@ 6 Conclusion
We introduce MIRAGE, a high-fidelity benchmark for evaluating vision-language models (VLMs) in expert-level agricultural consultations. MIRAGE addresses key limitations of existing benchmarks by incorporating real-world, multimodal, and context-rich interactions grounded in domain-specific decision-making. Despite these strengths, MIRAGE has a few limitations. It does not yet simulate interactive dialogue with real users or user simulators, limiting evaluation of adaptation and dialogue flow over time. Visual follow-ups are also not modeled; user turns beyond the initial user turn are assumed to be text-only. Looking ahead, MIRAGE opens several promising directions for advancing multimodal reasoning and conversational capabilities of models in knowledge-intensive domains. Improving performance on unseen entities will require better open-world generalization strategies, such as retrieval-augmented generation, compositional reasoning, or domain-adaptive pretraining. Limited gains from metadata highlight the need for explicit context modeling—including structured encodings of time, location, and integration with external knowledge bases capturing seasonal and biological priors. Persistent challenges in clarify-or-respond decision-making point to a need for models that can infer latent user goals and reason about missing information. We hope this benchmark enables the development of more context-sensitive, knowledge-intensive VLMs. MIRAGE is publicly released to support the development of vision-language systems that go beyond narrow question answering, toward models capable of engaging in natural interactions that involve ambiguity, a need for clarification, visual understanding, and decision-making when critical context is implicit or missing, addressing common challenges in real-world expert consultations.
7 Limitations

### v2

@@ 1 Introduction
Advances in large vision–language models (LVLMs) have significantly improved AI’s ability to interpret images and generate natural‐language responses. However, existing benchmarks predominantly focus on short-form visual question answering [yue2024mmmu, gauba2025agmmu, antol2015vqa, schwenk2022okvqa, zhang2024b], captioning [singla2024pixels, schuhmann2021laion], or grounded generation under constrained contexts [garcia2022improving]. These tasks fall short of capturing the interactive, decision-centric nature of real-world expert consultations, where users often present open-ended, ambiguous, and visually grounded queries. In knowledge-intensive domains such as agriculture, medicine, and engineering, expert consultations inherently span multiple modalities [salvi2024multi] and success hinges not just on perception or language fluency, but on the ability to reason causally, handle missing context, and make interaction-level decisions. This complex interplay between multimodal understanding and contextual reasoning in professional settings represents a significant gap in the current LVLM evaluation frameworks. LLMs and LVLMs are increasingly used in domains like healthcare, law, and plant care [tomar2024ecosage, wang2024agri, li2023llava, zhang2023huatuogpt, chen2024towards, ren2024healthcare, wu2024knowledge], yet current benchmarks underrepresent knowledge-intensive scenarios involving ambiguous, multimodal queries [gauba2025agmmu]. Agriculture illustrates this gap: farmers often seek image-based, expert-level guidance during critical events, but inaccurate or unsupported model outputs [zhai2023halle, liu2024survey, yin2024woodpecker] can lead to serious consequences. This highlights the need for rigorous evaluation frameworks that assess VLMs in more realistic settings.
@@ 2.1 Overview of MIRAGE
MIRAGE comprises two components: MMST, with a total corpus of over 29,000 single-turn expert-user interactions, and MMMT, a multi-turn corpus of 6,306 dialogues ( 12,000 turns) requiring clarify-or-respond decisions. Figure 1 provides a visual overview of the evaluation benchmarks derived from this corpus. Tasks span identification, diagnostic reasoning, and issue management guidance across seven agronomic categories (Table 4). The MMST evaluation benchmark includes Standard (8,184 interactions) and Contextual (3,934 interactions) subsets (Section 3, Appendix C), while the MMMT benchmark contains 861 dialogues. A fine-grained evaluation framework assesses both accuracy and utility. With over 7,600 biological entities in the full corpus (6,856 in the benchmark set), including thousands of unseen species in the evaluation set, MIRAGE introduces a challenging open-world generalization setting for LVLMs. Refer Appendix C for more details.
@@
| iNat21 [van2018inaturalist] | CLS | ✗ | ✓ | ✓ | – | – | – |
| TreeOfLife [stevens2024bioclip] | CLS | ✓ | ✓ | ✓ | – | – | – |
| SimpleQA [wei2024benchmarking] | OEQ | ✗ | ✗ | ✗ | ✓ | – | – |
| MMMU [yue2024mmmu] | MCQ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ |
| AgMMU [gauba2025agmmu] | MCQ+OEQ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| CROP [zhang2024empowering] | LFQ/Conv | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@ 2.2 Benchmark Comparison and Positioning
MIRAGE builds on a growing body of benchmarks in agriculture and multimodal reasoning, yet fills a critical gap left by prior efforts. Datasets like MMMU [yue2024mmmu] focus on generalist multimodal reasoning but rely on constrained multiple-choice formats. AgMMU [gauba2025agmmu] targets agriculture, yet uses synthetic short-form MCQs that are not representative of real expert-user interactions. Domain-specific datasets such as TreeOfLife [stevens2024bioclip] and iNat21 [van2018inaturalist] emphasize fine-grained image classification but operate in closed-world, non-interactive settings. The CROP benchmark [zhang2024empowering] introduces multi-turn crop science QA, but is limited to two crops, lacks visual input, and is fully text-based. In contrast, MIRAGE is the only benchmark constructed from large-scale, real-world agricultural consultations and supports both multimodal single-turn and multi-turn tasks grounded in naturally underspecified user queries. As summarized in Table 1, MIRAGE includes contextual question answering, real user-submitted images with varied quality and lighting, and domain-expert-authored responses as ground truth. It spans diverse query types—including identification, causal reasoning, and management—and introduces goal-state modeling, clarify-or-respond decision policies, and a multidimensional evaluation framework that moves beyond correctness to assess identification accuracy, causal justification, response quality, and diagnostic parsimony. These features make MIRAGE a uniquely realistic and challenging testbed for evaluating LVLMs in real-world consultation scenarios.
@@ 3 MIRAGE-MMST: Multimodal Singleturn Benchmark
MIRAGE-MMST is a benchmark designed to assess expert-level, single-turn reasoning in multimodal agricultural consultations. The task setup is similar to a Long-form VQA task [huh2024long]. Each instance consists of a natural language question paired with one or more user-provided images and associated metadata (e.g., timestamp, location). Each instance consists of a natural language question , an associated image set , and metadata . Formally, a single-turn instance is represented as a triplet , and the model must generate a structured response , where denotes identified entities (e.g., crop, pest, disease), is a causal explanation, and is a management recommendation, if requested. The task evaluates the model’s ability to reason causally about visual symptoms, identify relevant agronomic entities, and, when prompted, generate detailed management recommendations grounded in the observed evidence.
To support varying levels of difficulty and contextual grounding, MIRAGE-MMST is divided into two subsets: a Standard subset, consisting of self-contained questions that can be answered using only the provided text and image, and a Contextual subset, where successful interpretation depends on implicit information such as time, location, or agricultural context not present in the input. The Standard subset further includes two task types—MMST-ID (Identification), which focuses on visual reasoning and recognizing entities, and MMST-MG (Management), which involves reasoning and generating recommendation-based responses. In contrast, the Contextual subset contains queries with latent information gaps and elliptical language, requiring models to reconstruct missing context using external priors. We first manually annotated a seed set of contextual examples and then adopted an automated classifier to separate the full dataset. See Appendix C.3 for details.
@@ 4.1 Dataset Curation
The goal of MIRAGE-MMMT is to model the decision-making process of an expert who must decide, based on the evolving multimodal context, whether to respond with actionable guidance or clarify by seeking more information. To simulate this, we extract segments of each dialogue such that the input context ends with a user utterance, and a later user message, typically written in response to an earlier expert query, is treated as a revealed fact. This allows us to reconstruct the decision point at which the expert, based on available context and images, would determine whether the current information is sufficient to respond or whether a clarification is necessary. We prompt a powerful LLM (e.g., GPT-4o) to generate a structured task instance consisting of dialogue context, referenced images, and metadata such as geographic location and topic. To ensure data safety, we perform automated PII sanitization while preserving domain relevance. (see Appendix G.2 for more details)
@@ 5.1 Model Selection
The LVLMs we evaluated can be divided into two groups: 1.) Proprietary Models: this group comprises SOTA GPT models (4.1, 4.1-mini, 4o, 4o-mini) [achiam2023gpt], Claude-3 (3.5v2 Sonnet, 3.7 Sonnet, 3 Haiku) [anthropic2024claude] available through API service. 2.) Open-Source Models: this group includes LLaMa-4 Scout-17B [metaai2025llama4], LLaVa (7B mistral, 7B qwen2) [liu2023visual], Qwen 2.5 VL models (3B, 7B, 32B, 72B) [githubQwen3Qwen3_Technical_ReportpdfMain], Gemma 3 models (4b, 12, 27b) [team2025gemma] and Intern VL3 (2B, 8B, 14B, 38B, 72B) [zhu2025internvl3]. 3.) Finetuned Models We perform multi-image multimodal fine-tuning of the Qwen 2.5 VL models on the MMST dataset. All experiments were conducted with NVIDIA A100-40GB/H200-141GB GPUs.
@@ 5.2 Evaluation Methodology
To facilitate consistent, interpretable, and reproducible evaluation of model outputs, we implement a structured LLM-as-a-Judge protocol. Rather than relying on single-model, single-pass assessments, we leverage an ensemble of reasoning-capable language models: DeepSeek-R1-Distilled [guo2025deepseek], Qwen3-32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi-4-Reasoning [abdin2025phi], as judges. Each candidate’s response is evaluated across three independent generations, with every generation scored by the full ensemble, resulting in nine total evaluations per sample. This design captures both linguistic variability across generations and scoring consistency across evaluators. To validate the fidelity of this framework, we quantify inter-rater agreement using Fleiss’ for categorical scoring and Kendall’s W for ordinal coherence, ensuring both statistical reliability and scoring consistency across the ensemble. Fleiss’ scores for the ID task (binary classification) generally fall within the 0.75–0.88 range, indicating “good” to “excellent” agreement by standard interpretation guidelines. Full details, scoring rubrics, and agreement statistics are provided in Appendix D.1.1. The prompt for evaluation is showed in Figure 14.
@@ 5.3 MIRAGE-MMST Results
To evaluate model outputs on the MMST-MG task, we adopt a multidimensional scoring framework that assesses both factual correctness and communication quality. Each model response is rated on a 0–4 scale across the following four dimensions the judges: 1.) Accuracy: which measures the factual alignment with the expert’s response. A perfect score (4) is awarded when the facts, terminologies, and recommendations fully align with the expert response. Relevance measures the ability of the model to stay on-topic with the user’s original query and avoid tangential information. Completeness assesses whether the model covers all key information provided by the expert, and Parsimony, inspired by Occam’s Razor; this score captures the conciseness, clarity, and actionability of the model response. We aggregate these scores in a 2:1:1:1 ratio. See Appendix D for more details.
@@ 5.4 Performance Comparison
| Acc (%) | Reason | Acc | Rel | Comp | Pars | W‑Sum | |
@@
Scores are averaged over the three judge models. Closed‑source rows are red/light‑red, open‑source rows blue/light‑blue. Bold purple numbers mark the overall best for each metric.
@@
Model Finetuning: We utilize the MMST training dataset to fine-tune Qwen 2.5 VL models (3B, 7B, 32B) and evaluate performance on both MMST-ID & MMST-MG tasks. Figure 2 illustrates the effects of LoRA-based fine-tuning on Qwen2.5-VL-3B model across four checkpoints vs the base model (base-Instruct). Our results indicate consistent improvements in seen entity accuracy on MMST-ID after fine-tuning, rising from 22.3% (base-Instruct) to a peak of 28.4% (epoch 6). Across model sizes, as seen in Figure 10, we see a consistent convergence behavior across epochs, with 32B achieving the strongest peak performance (37.6% at epoch 6) and reasoning score (3.04). However, models struggle to generalize, as accuracy on unseen entities remains stagnant. On the MMST-MG task, from the base-Instruct model to LoRA-ep6, we observe steady improvement across all four metrics as well. Relevance is the highest-scoring metric throughout, indicating the model’s ability to remain on-topic is strong even in zero-shot, and improves further with tuning. Parsimony also improves consistently, indicating models can learn concise delivery of recommendations.
@@ 5.6 Performance Comparison
We evaluate the model’s high‐level decision and its alignment with user goals in three complementary ways. First, Decision Accuracy captures how often the model’s chosen action, whether to ask a clarification question or to respond directly, matches the gold annotation; To ensure that when the model generates an utterance, it is doing so meaningfully, we measure Clarify Relevance: for every turn labeled <Clarify> or Respond Relevance for all turns labeled <Respond> using the LLM judges.
@@
We conduct experiments under Zero-Shot and COT [wei2022chain] prompt settings on MMMT shown in Table 3. In the zero-shot setting, GPT-4o achieves the highest Decision Accuracy (62.98%) and leads in both Clarify Goal Relevance (70.75%) and Respond Goal Relevance (77.07%), showing strong overall reasoning and generation alignment. Open models like LLaMA 4 Maverick and LLaMA 4 Scout perform competitively, reaching 53–49% decision accuracy. Introducing chain-of-thought (CoT) reasoning consistently improves performance across models. GPT-4o gains +2.5% in Decision Accuracy (to 65.5%) and improves goal relevance in both Clarify (+2%) and Respond (+1.5%) settings. Other models, especially Claude 3.7 and Qwen 72B, benefit more substantially (improvements in clarify goal relevance by +10.5%).
@@ 6 Conclusion
We introduce MIRAGE, a high-fidelity benchmark for evaluating vision-language models (VLMs) in expert-level agricultural consultations. MIRAGE addresses key limitations of existing benchmarks by incorporating real-world, multimodal, and context-rich interactions grounded in domain-specific decision-making. Despite these strengths, MIRAGE has a few limitations. It does not yet simulate interactive dialogue with real users or user simulators, limiting evaluation of adaptation and dialogue flow over time. Visual follow-ups are also not modeled; user turns beyond the initial user turn are assumed to be text-only. Looking ahead, MIRAGE opens several promising directions for advancing multimodal reasoning and conversational capabilities of models in knowledge-intensive domains. Improving performance on unseen entities will require better open-world generalization strategies, such as retrieval-augmented generation, compositional reasoning, or domain-adaptive pretraining. Limited gains from metadata highlight the need for explicit context modeling, including structured encodings of time, location, and integration with external knowledge bases capturing seasonal and biological priors. Persistent challenges in clarify-or-respond decision-making point to a need for models that can infer latent user goals and reason about missing information. We hope this benchmark enables the development of more context-sensitive, knowledge-intensive VLMs. MIRAGE is publicly released to support the development of vision-language systems that go beyond narrow question answering, toward models capable of engaging in natural interactions that involve ambiguity, a need for clarification, visual understanding, and decision-making when critical context is implicit or missing, addressing common challenges in real-world expert consultations.
7 Limitations & Future Work